Americans combined drive about 4.0 x 109 miles per day and their vehicles get an average of 20 miles per gallon of fuel used, which is the hydrocarbon octane (C8H18).

a) write a balance equation for the oxidation of octane

b) then for each 1kg of gasoline that is burned, calculate how many kg of carbon dioxide are produced.

c) how mnay kilograms of CO2 are emitted into the atmosphere each day by cars in the U.S? The density of gasoline fuel is about 0.925g/ml (convert to Kg/gal)
Answer #1

(a )   C 8 H18 + 25/2 O 2 ---------> 8 CO 2 + 9 H 2O

( b )   From the above balanced equation:

              114 gm of Octane produces = 352 gm of CO2 is Produced

              1KG of Octane produces = 3.08KG of CO2 is Produced

( c)   Distance travelled = 4.0 *10 9 miles/day

         Average givenby the Vechile = 20 miles/ gallon

       Density of gasolinefuel  is = 0.925 g/ml  

                                                = 4.20 Kg/ gallon

    Total amount of fuel used perday = 4.0 * 10 9   miles/d ay / 20miles/gallon * 4.20 Kg/ gallon

                                                        = 8.40 * 10 8 Kg / day Used by thevechiles

       We know that 1 Kgof fuel produces = 3.08 Kg of CO2  

            8.40 * 10 8 Kg of fuel produces 25.872 * 10 8 Kg of CO2